# Accelerated ADMM: Automated Parameter Tuning and Improved Linear Convergence

### v1

@@ Abstract
This work studies the linear convergence of an accelerated scheme of the Alternating Direction Method of Multipliers (ADMM) for strongly convex and Lipschitz-smooth problems. We use the methodology of expressing the accelerated ADMM as a Lur’e system, i.e., an interconnection of a linear dynamical system in feedback with a slope-restricted operator, and we use Integral Quadratic Constraints to establish linear convergence. In addition, we propose several parameter tuning heuristics and their impact on the convergence rate through numerical analyses. Our new bounds show significantly improved linear convergence rates compared to the vanilla algorithm and previous proposed accelerated variants, which is also empirically validated on a LASSO regression benchmark.
@@ 1 Introduction
The Alternating Direction Method of Multipliers (ADMM) (Glowinski and Marroco, 1975; Boyd et al., 2011) has emerged as a benchmark algorithm among first-order methods for solving convex composite optimization problems. Its ability to decompose the original problem into subproblems that can be solved in parallel made it especially popular in computation-heavy domains as signal processing, machine learning or distributed control.
It is long established that ADMM enjoys provable convergence guarantees in the convex setting. In particular, Deng and Yin (2016) showed that for general convex problems ADMM achieves $\mathcal{O}(1/k)$ convergence rate, while for the strongly convex and smooth case it attains a linear convergence rate. An alternative proof of linear convergence was provided by Nishihara et al. (2015), who applied the Integral Quadratic Constraints (IQC) framework (Megretski and Rantzer, 1997; Lessard et al., 2016; Michalowsky et al., 2021) to numerically derive convergence rates and heuristics for parameter tuning. The proof frames the optimization algorithm as a Lur’e system (a feedback interconnection of an LTI system and monotone operators), and casts the convergence analysis as a stability problem. This framework offers the main advantage that also the analysis of altered schemes, such as over-relaxed variants, can be seamlessly carried out in an automated fashion as long as they can be framed in the Lur’e setting.
Also accelerated variants of ADMM (A-ADMM) have been proposed and analyzed in recent years. Goldstein et al. (2014) showed that Nesterov acceleration leads to a convergence rate of $\mathcal{O}(1/k^{2})$ in the strongly convex settings. Wang et al. (2021) provided linear convergence guarantees when Anderson acceleration is applied. Patrinos et al. (2014); Pejcic and Jones (2016) proved linear convergence of accelerated Douglas–Rachford splitting and ADMM for strongly convex quadratic problems, showing that the optimal rate matches that of Nesterov’s fast gradient method. However, an analysis of the convergence rates of A-ADMM for general strongly convex and smooth problems is still unaddressed, as well as the question of whether other known rates from classical accelerated gradient methods can be matched.
@@
$\min f(x)$ Quadr. Nesterov Method $\rho_{1}$ Nesterov (2004)
$\min f(x)$ $S(m,L)$ Nesterov Method $\rho_{2}$ Safavi et al. (2018)
$\min f(x)$ $S(m,L)$ Triple Momentum $\rho_{1}$ Van Scoy et al. (2018)
(1) Quadr. A-ADMM (NM) $\rho_{1}$ Patrinos et al. (2014); Pejcic and Jones (2016)
@@
In this work, we extend the IQC-based analysis of vanilla ADMM (Nishihara et al., 2015) to the accelerated case, when one of the objectives is strongly convex and smooth. By casting the A-ADMM algorithm as a Lur’e system, we propose a semi-definite program (SDP) that provides a numerical tool to theoretically verify worst-case convergence rates of accelerated schemes. Using dynamic O’Shea-Zames-Falb (OZF) IQCs (Boczar et al., 2015), we prove that A-ADMM achieves tighter worst-case convergence bounds than vanilla ADMM. Moreover, we propose a systematic parameter tuning procedure based on this framework and demonstrate that Nesterov- and Triple-Momentum-inspired parameter selections attain rates close to those of accelerated first-order methods (cf. Table 1). Finally, we identify through grid search a new A-ADMM configuration with the fastest certified convergence rate, which also exhibits the best empirical performance in our numerical experiments. Our theoretical results are validated on a LASSO regression case study, showing improved convergence speed of all our schemes compared to existing benchmark algorithms.
@@ 2.1 Accelerated ADMM
where $f:\mathbb{R}^{p}\to\mathbb{R}$, $f\in S_{p}(m,L)$ and $g:\mathbb{R}^{q}\to\mathbb{R}$, $g\in S_{q}(0,\infty)$. The matrices $A\in\mathbb{R}^{p\times p}$, $B\in\mathbb{R}^{p\times q}$, and $c\in\mathbb{R}^{p}$ define linear equality constraints coupling the variables $x$ and $z$, where we particularly assume that $A$ is invertible and that $B$ has full column rank. Having a strongly convex and smooth component in (1) is common in many practical applications (Deng and Yin, 2016), while restricting $A$ to be square and invertible is slightly restrictive, but also done in Nishihara et al. (2015). We note that still many practical problem instances satisfy this structure, including consensus problems and various distributed optimization tasks (Notarstefano et al., 2019).
To address problem (1), we use the over-relaxed ADMM algorithm (Boyd et al., 2011)
@@
$z_{k+1}$ $=\arg\min_{z}\;g(z)+\tfrac{1}{2\nu_{1}}\,\|\alpha Ax_{k+1}-(1-\alpha)Bz_{k}$ (2b)
@@
$\lambda_{k+1}$ $=\alpha Ax_{k+1}-(1-\alpha)Bz_{k}+Bz_{k+1}-\alpha c+\hat{\lambda}_{k}$ (2c)
@@
as in Goldstein et al. (2014); Pejcic and Jones (2016). Here, $\nu_{1}>0$ is a step size, $\nu_{2}\geq 0$ a momentum parameter and $\alpha\geq 0$ the so-called over-relaxation parameter. We recover the nominal A-ADMM by setting $\alpha=1$, and the vanilla, non-accelerated ADMM with $\nu_{2}=0$. Both $\alpha$ and $\nu_{2}$ have been shown to improve the convergence speed for suitable choices (Eckstein and Bertsekas, 1992; Patrinos et al., 2014). The over-relaxation parameter $\alpha$ is chosen in the interval $(0,2]$, while $\nu_{2}$ is often selected according to Nesterov’s schemes (Nesterov, 2004), e.g., $\nu_{2}=(\sqrt{L}-\sqrt{m})/(\sqrt{L}+\sqrt{m})$ (Patrinos et al., 2014), or as an iteration dependent quantity (Goldstein et al., 2014). To explore alternative parameter selection strategies we employ the IQC framework.
@@ 2.2 Integral Quadratic Constraints for Convex Functions
Algorithm (2)-(3) will be analyzed as a dynamical system converging to points satisfying the first-order optimality conditions. We will make use of the well-known fact that (sub)gradients of (strongly) convex functions are monotone and, thus, their input/output relation satisfy IQCs for slope-restricted operators (Lessard et al., 2016). To be more precise, they have been shown to fulfill so-called $\rho$-hard O’Shea-Zames–Falb IQCs.
Proposition 1 ($\rho$-OZF IQC, Boczar et al. (2015))
@@
The filter coefficients $h^{f}$, $h^{g}$ will serve as degree-of-freedom in the resulting SDP, subject to the convex constraint (4). In theory, the filter dimension $n_{\mathrm{OZF}}$ can be infinite, but is chosen finite in practice to yield finite-dimensional state-space realizations. A choice of $n_{\mathrm{OZF}}=0$ leads to static filters $\Psi_{f},\Psi_{g}$ and pointwise satisfaction of the inequalities (6), which in literature is often referred to as sector IQC (Lessard et al., 2016).
@@ 2.3 Stability Analysis of First-Order Algorithms
The proof comes as a straightforward extension of (Lessard et al., 2016, Theorem. 4) to the two-operator case, which particularly exploits the positivity constraint (6). Eq. (11) is independent of the problem dimension $p$, and can be used to determine the minimum worst-case convergence rate $\rho$ via bisection. Note that Theorem 2 recovers (Nishihara et al., 2015, Theorem. 6) for $n_{\mathrm{OZF}}=0$.
@@ 3 A-ADMM as a Dynamical System
Building on the previous derivations, we now formulate the A-ADMM (2)-(3) as a Lur’e system (7). In line with Nishihara et al. (2015), we introduce the coordinate change $r_{k}=Ax_{k}$, $s_{k}=Bz_{k}$. Moreover, define
@@
In contrast to Nishihara et al. (2015), we deliberately retain the ADMM step-size $\nu_{1}$ in the state-space description, instead of putting it in the sector bounds. The momentum parameter $\nu_{2}$ becomes an additional degree-of-freedom.
@@ 4 Parameter Selection And Additional Modifications
In unconstrained first-order methods, accelerated algorithms are parametrized by a step-size, a damping-parameter, and an extrapolation-parameter, where each quantity has provable optimal tunings (Van Scoy et al., 2018). For ADMM in contrast, convergence proofs and parameter selection of $\nu_{1},\nu_{2},\alpha$ have been traditionally difficult. In this section we explore different parameter selection heuristics and show the influence on the convergence rates $\rho$ that are obtained from Theorem 2. As initial heuristics, we adopt the same parameter tuning schemes of accelerated gradient methods. Next, we perform a grid-search to identify potentially superior configurations.
@@ 4.1 Nesterov and Triple Momentum Parameters
As heuristics, we use here the step-size and damping parametrization of the Nesterov method (NM) (Nesterov, 2004) and the Triple Momentum algorithm (TM) (Van Scoy et al., 2018), which is known to attain the fastest possible convergence rate for accelerated gradient methods. The parameter selections are summarized in Table 2. We plot the resulting $\rho$ obtained by Theorem 2 as a function of the condition number $\kappa=\hat{L}/\hat{m}$. We increase the O’Shea-Zames-Falb order until $\rho$ does not improve anymore, giving $n_{\mathrm{OZF}}=6$ in all experiments [^1] Code for all experiments can be found on https://github.com/col-tasas/2025-accelerated-admm . Fig. 1 summarizes the results for both settings, in comparison to the rates obtained from vanilla ADMM. We note that the rates of vanilla ADMM have been shown to be tight (Nishihara et al., 2015), while the curves are upper bounds.
For the NM parameter set, observe in Fig. 1(a) that the convergence rates substantially outperform vanilla ADMM, certifying better linear convergence across all condition numbers. Moreover, we observe that the rates asymptotically approach the optimal rate of Nesterov’s accelerated method, given by $\rho={\scriptsize\sqrt{1-\frac{\sqrt{2\kappa-1}}{\kappa}}}$ (Safavi et al., 2018). For the TM parameter set, observe in Fig. 1(b) that the convergence rates outperform vanilla ADMM in low condition regions, but from $\kappa\approx 18$ onwards, the rates degrade and eventually become infeasible. Thus, no statement about the convergence for higher condition ratios can be done. We now propose a modification to the ADMM update to alleviate this issue.
@@ 4.2 $\lambda$-Damping
The results of A-ADMM with TM parameters and additional $\lambda$-damping with $d=0.1$ are shown in Fig. 1(b). We observe that convergence can now be guaranteed for all condition ratios, and as well that the rates are better than for vanilla ADMM. Moreover, we observe analogous to the NM case that the convergence rate curve approaches the optimal rate of Triple Momentum, given by $\rho=1-\frac{1}{\sqrt{\kappa}}$. This suggests that the additional damping of the dual variable extends the range of provable convergence for TM A-ADMM.
@@ 4.3 Grid Search Results
In addition to the already promising TM and NM parameter selections, we perform a grid search to identify potentially improved configurations of $(\nu_{1},\nu_{2})$. Our search is set up as a $20\times 20$ grid centered around the TM parameters. As commonly done in strongly-convex and smooth optimization problems, we parametrize the parameters as a function of the condition ratio $\kappa$. For each $\kappa$ we select $(\nu_{1},\nu_{2})$ as the pair for which the minimal $\rho$ can be obtained with Theorem 2 via bisection. The procedure is repeated for the over-relaxed A-ADMM (OR-A-ADMM), with an additional grid over the relaxation parameter $\alpha$ in the range $(0,4]$. Note that even though the standard theory would prevent from exploring the range $(2,4)$, we can still perform this interval extension and check for linear convergence by the feasibility of (11).
@@
Fig. 2 shows the grid search results for $\nu_{1}$ and $\nu_{2}$. We found that the stepsize parameter $\nu_{1}$ found via grid search exactly coincides with the TM value. In contrast, the step-size parameter $\nu_{2}$ exhibits a different behavior. The same holds for the over-relaxed case. Applying symbolic regression (Cranmer, 2023) reveals that $\nu_{2}$ can be approximated by
@@
Fig. 3 summarizes the convergence rates obtained from the grid search (GS) parameters, in comparison to the vanilla ADMM method and the accelerated versions with NM and TM parameters. By design, the A-ADMM with GS-parameters gives the best linear convergence rate of A-ADMM among all settings. Fig. 4 complements the results with the over-relaxed variants. We observe that the over-relaxed variants constantly outperform the non-over-relaxed ones; and that the accelerated variants constantly outperform the non-accelerated variants. Moreover, A-ADMM also outperforms the vanilla over-relaxed ADMM.

### v2

@@ Abstract
This work studies the linear convergence of an accelerated scheme of the Alternating Direction Method of Multipliers (ADMM) for strongly convex and Lipschitz-smooth problems. We use the methodology of expressing the accelerated ADMM as a Lur’e system, i.e., an interconnection of a linear dynamical system in feedback with a slope-restricted operator, and we use Integral Quadratic Constraints to establish linear convergence. In addition, we propose several parameter tuning heuristics and their impact on the convergence rate through numerical analyses. Our new bounds show improved linear convergence rates compared to the vanilla algorithm and previous proposed accelerated variants, which is also empirically validated on a LASSO regression benchmark.
@@ 1 Introduction
The Alternating Direction Method of Multipliers (ADMM) (Glowinski1975; boydAdmm) has emerged as a benchmark algorithm among first-order methods for solving convex composite optimization problems. Its ability to decompose the original problem into subproblems that can be solved in parallel made it especially popular in computation-heavy domains as signal processing, machine learning or distributed control.
It is long established that ADMM enjoys provable convergence guarantees in the convex setting. DengYin showed that for general convex problems ADMM achieves a $\mathcal{O}(1/k)$ convergence rate, while for the strongly convex and smooth case it attains a linear convergence rate, with optimally tuned parameters given in pontus_boyd. An alternative proof of linear convergence was presented by NishiharaADMM, who applied the Integral Quadratic Constraints (IQC) framework (Megretski1997; LessardRechtPackard; SchererEbenbauer) to numerically derive convergence rates and tune parameters. The proof frames the optimization algorithm as a Lur’e system (a feedback interconnection of an LTI system and monotone operators), and casts the convergence analysis as a stability problem. This framework offers the main advantage that also the analysis of altered schemes, such as over-relaxed variants, can be seamlessly carried out in an automated fashion as long as they can be framed in the Lur’e setting.
Also accelerated variants of ADMM (A-ADMM) have been proposed and analyzed in recent years. FastADMM showed that Nesterov acceleration leads to a convergence rate of $\mathcal{O}(1/k^{2})$ in the strongly convex settings. Sterck provided linear convergence guarantees when Anderson acceleration is applied. PatrinosStellaBemporad; PejcicJones proved linear convergence of accelerated Douglas–Rachford splitting and ADMM for strongly convex quadratic problems, showing that the optimal rate matches that of Nesterov’s fast gradient method. However, an analysis of the convergence rates of A-ADMM for general strongly convex and smooth problems is still unaddressed, as well as the question of whether other known rates from classical accelerated gradient methods can be matched.
@@
$\min f(x)$ Quadr. Nesterov Method $\rho_{1}$ Nesterov2004
$\min f(x)$ $S(m,L)$ Nesterov Method $\rho_{2}$ safavi
$\min f(x)$ $S(m,L)$ Triple Momentum $\rho_{1}$ Freeman2028
(1) Quadr. A-ADMM (NM) $\rho_{1}$ PatrinosStellaBemporad; PejcicJones
@@
In this work, we extend the IQC-based analysis of vanilla ADMM (NishiharaADMM) to the accelerated case, when one of the objectives is strongly convex and smooth. By casting the A-ADMM algorithm as a Lur’e system, we propose a semi-definite program (SDP) that provides a numerical tool to theoretically verify worst-case convergence rates of accelerated schemes. Using dynamic O’Shea-Zames-Falb (OZF) IQCs (BoczarLessardRecht), we propose a systematic parameter tuning procedure and demonstrate that Nesterov- and Triple-Momentum-inspired parameter selections attain rates close to those of accelerated first-order methods (cf. Table 1). Finally, we identify through grid search a new A-ADMM configuration with the fastest certified convergence rate, outperforming all vanilla ADMM methods. Our theoretical results are validated on a LASSO regression case study, demonstrating consistently improved convergence speed of our schemes compared to existing benchmark algorithms.
@@ 2.1 Accelerated ADMM
where $f:\mathbb{R}^{p}\to\mathbb{R}$, $f\in S_{p}(m,L)$ and $g:\mathbb{R}^{q}\to\mathbb{R}$, $g\in S_{q}(0,\infty)$. The matrices $A\in\mathbb{R}^{p\times p}$, $B\in\mathbb{R}^{p\times q}$, and $c\in\mathbb{R}^{p}$ define linear equality constraints coupling the variables $x$ and $z$, where we particularly assume that $A$ is invertible and that $B$ has full column rank. Having a strongly convex and smooth component in (1) is common in many practical applications (DengYin), while restricting $A$ to be square and invertible is slightly restrictive, but also done in NishiharaADMM. We note that still many practical problem instances satisfy this structure, including consensus problems and various distributed optimization tasks (NotarstefanoNotarnicolaCamisa).
To address problem (1), we use the over-relaxed ADMM algorithm (boydAdmm)
@@
$z_{k+1}$ $=\arg\min_{z}\;g(z)+\tfrac{1}{2\nu_{1}}\,\|\alpha Ax_{k+1}-(1-\alpha)B\hat{z}_{k}$ (2b)
@@
$\lambda_{k+1}$ $=\alpha Ax_{k+1}-(1-\alpha)B\hat{z}_{k}+Bz_{k+1}-\alpha c+\hat{\lambda}_{k}$ (2c)
@@
as in FastADMM; PejcicJones. Here, $\nu_{1}>0$ is a step size, $\nu_{2}\geq 0$ a momentum parameter and $\alpha\geq 0$ the so-called over-relaxation parameter. We recover the nominal A-ADMM by setting $\alpha=1$, and the vanilla, non-accelerated ADMM with $\nu_{2}=0$. Both $\alpha$ and $\nu_{2}$ have been shown to improve the convergence speed for suitable choices (Eckstein1992; PatrinosStellaBemporad). The over-relaxation parameter $\alpha$ is chosen in the interval $(0,2]$, while $\nu_{2}$ is often selected according to Nesterov’s schemes (Nesterov2004), e.g., $\nu_{2}=(\sqrt{L}-\sqrt{m})/(\sqrt{L}+\sqrt{m})$ (PatrinosStellaBemporad), or as an iteration dependent quantity (FastADMM). To explore alternative parameter selection strategies we employ the IQC framework.
@@ 2.2 Integral Quadratic Constraints for Convex Functions
Algorithm (2)-(3) will be analyzed as a dynamical system converging to points satisfying the first-order optimality conditions. We will make use of the well-known fact that (sub)gradients of (strongly) convex functions are monotone and, thus, their input/output relation satisfy IQCs for slope-restricted operators (LessardRechtPackard). To be more precise, they have been shown to fulfill so-called $\rho$-hard O’Shea-Zames–Falb IQCs.
Proposition 1 ($\rho$-OZF IQC, BoczarLessardRecht)
@@
The filter coefficients $h^{f}$, $h^{g}$ will serve as degree-of-freedom in the resulting SDP, subject to the convex constraint (4). In theory, the filter dimension $n_{\mathrm{OZF}}$ can be infinite, but is chosen finite in practice to yield finite-dimensional state-space realizations. A choice of $n_{\mathrm{OZF}}=0$ leads to static filters $\Psi_{f},\Psi_{g}$ and pointwise satisfaction of the inequalities (6), which in literature is often referred to as sector IQC (LessardRechtPackard).
@@ 2.3 Stability Analysis of First-Order Algorithms
The proof comes as a straightforward extension of (LessardRechtPackard, Theorem. 4) to the two-operator case, which particularly exploits the positivity constraint (6). Eq. (11) is independent of the problem dimension $p$, and can be used to determine the minimum worst-case convergence rate $\rho$ via bisection. Note that Theorem 2 recovers (NishiharaADMM, Theorem. 6) for $n_{\mathrm{OZF}}=0$.
@@ 3 A-ADMM as a Dynamical System
Building on the previous derivations, we now formulate the A-ADMM (2)-(3) as a Lur’e system (7). In line with NishiharaADMM, we introduce the coordinate change $r_{k}=Ax_{k}$, $s_{k}=Bz_{k}$. Moreover, define
@@
In contrast to NishiharaADMM, we deliberately retain the ADMM step-size $\nu_{1}$ in the state-space description, instead of putting it in the sector bounds. The momentum parameter $\nu_{2}$ becomes an additional degree-of-freedom.
@@ 4 Parameter Selection And Additional Modifications
In unconstrained first-order methods, accelerated algorithms are parametrized by a step-size, a damping-parameter, and an extrapolation-parameter, where each quantity has provable optimal tunings (Freeman2028). For ADMM in contrast, convergence proofs and parameter selection of $\nu_{1},\nu_{2},\alpha$ have been traditionally difficult. In this section we explore different parameter selection heuristics and show the influence on the convergence rates $\rho$ that are obtained from Theorem 2. As initial heuristics, we adopt the same parameter tuning schemes of accelerated gradient methods. Next, we perform a grid-search to identify potentially superior configurations.
@@ 4.1 Nesterov and Triple Momentum Parameters
We investigate the use of the parameterization of the Nesterov method (NM) (Nesterov2004) and the Triple Momentum algorithm (TM) (Freeman2028), which is known to attain the fastest possible convergence rate for accelerated gradient methods. The parameter selections are summarized in Table 2. We note that it had already been proposed to tune A-ADMM based on the NM parameter set (PejcicJones). We plot the resulting $\rho$ obtained by Theorem 2 as a function of the condition number $\kappa=\hat{L}/\hat{m}$. We increase the O’Shea-Zames-Falb order until $\rho$ does not improve anymore, giving $n_{\mathrm{OZF}}=6$ in all experiments [^1] Code for all experiments can be found on https://github.com/col-tasas/2025-accelerated-admm .
For the NM parameter set, observe in Fig. 1(a) that the convergence rates indeed approach the optimal rate of Nesterov’s accelerated method, given by $\rho={\scriptsize\sqrt{1-\frac{\sqrt{2\kappa-1}}{\kappa}}}$ (safavi). For the TM parameter set, observe in Fig. 1(b) that the convergence rates also approach the optimal rate of Triple Momentum, given by $\rho=1-\frac{1}{\sqrt{\kappa}}$, However, from $\kappa\approx 18$ onwards, the rates degrade and eventually become larger than one. Thus, no statement about the convergence for higher condition ratios can be done. We now propose a modification to the ADMM update to alleviate this issue.
@@ 4.2 $\lambda$-Damping
The results of A-ADMM with TM parameters and additional $\lambda$-damping with $d=0.1$ are shown in Fig. 1(b). We observe that convergence can now be guaranteed for all condition ratios, and moreover, we observe analogous to the NM case that the convergence rate curve approaches the optimal rate of Triple Momentum, given by $\rho=1-\frac{1}{\sqrt{\kappa}}$. This suggests that the additional damping of the dual variable extends the range of provable convergence for TM A-ADMM, and that the strong connection between A-ADMM and accelerated first-order methods is also preserved in the strongly-convex and smooth setting.
@@ 4.3 Grid Search Results
We now perform a grid search to identify potentially improved configurations of $(\nu_{1},\nu_{2})$. Our search is set up as a $20\times 20$ grid centered around the TM parameters. As commonly done in strongly-convex and smooth optimization problems, we parametrize the parameters as a function of the condition ratio $\kappa$. For each $\kappa$ we select $(\nu_{1},\nu_{2})$ as the pair for which the minimal $\rho$ can be obtained with Theorem 2 via bisection. The procedure is repeated for the over-relaxed A-ADMM (OR-A-ADMM), with an additional grid over the relaxation parameter $\alpha$ in the range $(0,4]$. Note that even though the standard theory would prevent from exploring the range $(2,4)$, we can still perform this interval extension and check for linear convergence by the feasibility of (11).
@@
Fig. 2 shows the grid search results for $\nu_{1}$ and $\nu_{2}$. We found that the stepsize parameter $\nu_{1}$ found via grid search exactly coincides with the TM value. In contrast, the step-size parameter $\nu_{2}$ exhibits a different behavior. The same holds for the over-relaxed case. Applying symbolic regression (pysr) reveals that $\nu_{2}$ can be approximated by
@@
Fig. 3 summarizes the convergence rates obtained from the grid search (GS) parameters, in comparison to the vanilla ADMM method and the accelerated versions with NM and TM parameters. By design, the A-ADMM with GS-parameters gives the best linear convergence rate of among all settings. Fig. 4 complements the results with the over-relaxed variants. We observe that the over-relaxed variants constantly outperform the non-over-relaxed ones; and that the accelerated variants constantly outperform the non-accelerated variants.